what is the glucose that can be produced from a photosynthesis reaction that occurs using 10 mol CO2 using the equation 6CO2 + 6H2O --> C6H12O6 + 6O2?


1
Expert's answer
2021-05-24T07:17:02-0400

v(C6H12O6)=v/6(CO2)=10mol/6=1.67mol (by the equation)

m(C6H12O6)=v(C6H12O6)*M(C6H12O6)

M(C6H12O6)=6*12g/mol+12*1g/mol+6*16g/mol=180g/mol (аccording to the periodic table)

m(C6H12O6)=1.67mol*180g/mol=300g